What average force is required to stop a 1900 kg car in 8.0 s if the car is traveling at 100 km/h ?

Express your answer to two significant figures and include the appropriate units. Enter positive value if the direction of the force is in the direction of the initial velocity and negative value if the direction of the force is in the direction opposite to the initial velocity.
Answer #1

First of all, convert the speed , v = 100 km/h into meter/s.

So –

v = 100 km/h = (100 x 1000) / (60 x 60) m/s = 27.78 m/s

Now, find out the retardation of the car –

v’ = v + a*t

here, v’ = final speed of the car = 0

=> 0 = 27.78 + a*8

=> a = -27.78/8 = -3.47 m/s^2

The negative sign above shows the retardation of the car.

Now. mass of the car, m = 1900 kg

Therefore, average force required to stop the car, F = m*a = 1900*(-3.47) = - 6593.0 N

In two significant digits, our answer = -6600.0 N

Please note the negative sign which shows that the direction of application of the force is opposite to the direction of motion of the car.